OpenStudy (bohotness):

. Find the cube of 12. A. 36 B. 144 C. 1728 D. 20,736 help pelase

OpenStudy (bohotness):

@WhiteFangIII

OpenStudy (bohotness):

??

OpenStudy (anonymous):

idk i got something else. . but try c

OpenStudy (princeharryyy):

12*12*12 = 1728

OpenStudy (princeharryyy):

@bohotness

OpenStudy (anonymous):

so i was right ? @princeharryyy

OpenStudy (bohotness):

yeaa

OpenStudy (bohotness):

i throught its was b

OpenStudy (princeharryyy):

yes u were @TrapBoiTrynaEducaate u were right on y=sinx the other guy was messing without any reason

OpenStudy (princeharryyy):

no prob... now u got it correct hmmm @bohotness

OpenStudy (anonymous):

its fine i didnt care much lol let them be

OpenStudy (bohotness):

:)

OpenStudy (princeharryyy):

hmmm hehe :) idk he thought himself to b smarter...

OpenStudy (bohotness):

2.Find the volume of a cube with the given side length. 8 units A. 24 cubic units B. 64 cubic units C. 512 cubic units D. 4096 cubic units d??

OpenStudy (anonymous):

flutter u hahaha u didnt give me my medal

OpenStudy (princeharryyy):

no problem i give u that....))) lets not fight over here

OpenStudy (bohotness):

:P

OpenStudy (princeharryyy):

i gave him @bohotness

OpenStudy (bohotness):

ok

OpenStudy (princeharryyy):

he is nice too ... ))

OpenStudy (princeharryyy):

8*8*8 = 512 as volume of cube = (side)*(side)*(side)

OpenStudy (princeharryyy):

@bohotness

OpenStudy (bohotness):

so just time them three times??

OpenStudy (anonymous):

id listen to prince harry hes more skilled in these mathematics then me :P

OpenStudy (princeharryyy):

yes, but only in case of cube.

OpenStudy (bohotness):

okay

OpenStudy (princeharryyy):

:))))

OpenStudy (anonymous):